Issues which arose during the 7-18 June 2010 Science Run

LBC: FPIA using incorrect configuration file

Observers had a copy of the file lbcfpia.cfg in their home directory, the directory from which they were running dofpia. As a consequence, instead of using the latest (and default) configuration file in /home/lbcobs/LBCFPIA/lbcfpia/src/ which has the matrix and new Z22 zeropoints, they were using this older version. The Z22 zero points were incorrect. What alerted us to a problem was the Red guide star thumbnails, where the star was way out of focus - a donut. Dofpia prints the Z22 zeropoints in use, and we then realized that these were not the current values.

We need some way to insure that the correct version of the software and associated files are being run.

LBC: Image Quality problem at I-BESSEL (and other LBC-Red filters, V-BESSEL)

First image through I-BESSEL: there is coma on tech chip 1 - noticeable in the thumbnail. Resulting 150 sec image is good at center but degrades on the edges, especially on chip 1 (chip 4 not examined).

The filter focus offset is good, as an image taken on 5 June 2010 shows better image quality throughout.

Compare lbcr.20100605.084957.fits and lbcr.20100618.090641.fits or lbcr.20100618.092849.fits. On 20100605, dofpia sequence prior to good image was:
  -950      0      0      0      0    500   0    y  0.00  0.00      dofpia_sph3_adjustment                
  -278    204     12    103   -183     14  -3    y  0.77  0.18       /newdata/lbcr.20100605.083120.fits  r-SLOAN 
  -465   -158    -80   -224     61     28  -1    y  0.67  0.23       /newdata/lbcr.20100605.083248.fits  r-SLOAN 
   243   -212     43    278   -194    -66 -16    y  0.74  0.21       /newdata/lbcr.20100605.083415.fits  r-SLOAN 
  -163   -129    -75    -27    115     29 -11    y  0.70  0.12       /newdata/lbcr.20100605.083539.fits  r-SLOAN 
   608      0      0      0      0   -320   0    y  0.00  0.00      dofpia_sph3_adjustment                
On 20100618, dofpia sequence was:
  -950      0      0      0      0    500   0    y  0.00  0.00   dofpia_sph3_adjustment
   851    -35   -214    324    -53   -344   0    y  0.55  0.44    /newdata/lbcr.20100618.090211.fits  r-SLOAN
    58    132    253   -386     68    -33  -6    y  0.52  0.21    /newdata/lbcr.20100618.090333.fits  r-SLOAN
    -6    -70    -16    156      0    -21   9    y  0.45  0.11    /newdata/lbcr.20100618.090453.fits  r-SLOAN
   438      0      0      0      0   -230   0    y  0.00  0.00   dofpia_sph3_adjustment

Pupils from lbcr.20100605.083539 and lbcr.20100618.090453 look very similar. See attached picture. Mirror position looks similar for both dates/images:
ecl> hsel /d2/kuhn/LBCdata/20100618/lbcr.20100618.090641.fits[0] $I,mirrorx,mirrory,mirrorz,mirrorrx,mirrorry,z4,z5,z6,z7,z8,z11,z22
boolean expression governing selection (yes): 
/d2/kuhn/LBCdata/20100618/lbcr.20100618.090641.fits[0]  1.102   -0.801  0.743   -3.550  -19.629 438.61000       0.00000 0.00000 0.00000 0.00000 -230.85000      0.00000
ecl> hsel /d2/kuhn/LBCdata/20100605/sn09/lbcr.20100605.083801.fits[0] $I,mirrorx,mirrory,mirrorz,mirrorrx,mirrorry,z4,z5,z6,z7,z8,z11,z22
boolean expression governing selection (yes): 
/d2/kuhn/LBCdata/20100605/sn09/lbcr.20100605.083801.fits[0]     1.602   -1.235  0.948   -2.351  -23.504 608.35000       0.00000 0.00000 0.00000 0.00000 -320.18000      0.00000
To remedy this, observers started to apply Z4 corrections to DX. However, was automatic tech chip focussing removing these? Need to look at this in more detail.

Note added 20 June: The image quality and PSF pattern is similar on 20100605.083801 and 20100618.090641. I had not realized this earlier, but on 5 June there is a sign of coma on the edge of chip 1 and the guide star on lbcrtec.20100605.083811_1.fits is not round. It is not as aberrated as on lbcrtec.20100618.090652_1.fits but the same trend is there.

I do not see much evidence for this off-axis coma on 20100617.102717 (I-Bessel; FWHMscl=4.89), however I do see it on the sharpest I-Bessel image later that night, 20100617.104845 (FWHMscl = 3.93). The image from 18 June had FWHMscl = 3.36, and the one from 5 June was sharper.

LBC: non-zero offsets for first dither position are ignored

This was issuetrak #467. It was fixed, but since then changes have been made both to the way LBC commands presets & offsets and probably to the TCS also.

First dither offset is ignored. Images/programs affected:

20100613: Rigliaco_Serpens_U Rigliaco
20100614: WD_N6791_LONG_01 Piotto
20100615: WD_N6791_LONG_01 and WD_N6791_LONG_02 Piotto
20100616: Rigliaco_Serpens_U Rigliaco

LBC: Offsets are inaccurate

Checked the accuracy of offsets for an OB with large offsets. The target was at 19:20:53 37:48:18

Commanded offsets were:
1. lbcr.20100615.090257.fits 19:20:52.999 +37:46:18.02 -288 -288 (first offset, ignored, IT #467) 
2. lbcr.20100615.091845.fits 19:20:34.801 +37:41:38.93 -216 -279
3. lbcr.20100615.093439.fits 19:20:40.867 +37:41:47.98 -144 -270

Ran imexam to measure the (x,y) centers for 3 stars on chip 2.  

The telra, teldec reflect the correct offsets, but the differences in positions of 3 stars on
LBC images (concentrating only on CHIP2 and assuming 0.225 and 0.228 arc/pix scales) would
predict offsets different from those commanded by about 6-10 arcseconds. 

# dx12 dy12 dox12 doy12     dx13 dy13 dox13 doy13       dx23 dy23 dox23 doy23
#
# assuming scale = 0.225
# red
star1. 208.6 -276.6 216.0 279.0    135.5 -262.9 144.0 270.0    -73.1 13.6 -72.0 -9.0 
star2. 209.8 -278.2 216.0 279.0    135.2 -261.8 144.0 270.0    -74.6 16.4 -72.0 -9.0 
star3. 210.2 -275.5 216.0 279.0    134.6 -263.0 144.0 270.0    -75.5 12.6 -72.0 -9.0 
# dx12 dy12 dox12 doy12     dx13 dy13 dox13 doy13        dx23 dy23 dox23 doy23
#
# assuming scale 0.228 arcsec/pixel
# red camera
star1. 211.4 -280.3 216.0 279.0    137.3 -266.4 144.0 270.0    -74.1 13.8 -72.0 -9.0 
star2. 212.6 -281.9 216.0 279.0    137.0 -265.3 144.0 270.0    -75.6 16.6 -72.0 -9.0 
star3. 213.0 -279.2 216.0 279.0    136.4 -266.5 144.0 270.0    -76.5 12.7 -72.0 -9.0 

LBC: Must correct pointing and co-pointing for observations at EL>80 degrees

LBC Red images 092231, 092627 and 093027: Stellar PSF is elongated with a small blip at the top.
LBC Blue image 092240 stars are very elongated, 093034 also elongated.

Blue master and Red slave.

EL = 83 degrees.

We think that the Blue was not tracking well near zenith and the red mirror may have jumped in guiding to keep up.

LUCIFER: Wait for collimation status must be "true" and YES for LUCIFER data-taking to wait.

GCS: vibrations occur at low elevation (< 40 degrees)

Tracking a target from 57 to 38 degrees EL, we had to stop because we noticed that the guide star was extremely elongated - about 10:1 axis ratio. Called Torsten & Dave Thompson. There is a problem with vibrations and the guide star being on the edge of the control field.

LBC: SX collimation at start requires z4 and coma corrections.

At the start of the night, with T~11C, the first pupil images are too small. They require z4 ~ -15,000 nm. Typically they also require some coma (z7).

Network slow at 4am 18 June 2010.

LBC images took a long time to appear in newdata. During sky flats, this delay is critical.

-- OlgaKuhn - 16 Jun 2010

  • Comparison of pupils on 5 June and 18 June 2010.:
    compare_pupils.jpg

  • right side of chip 1 from lbcr.20100618.090641, I-Bessel:
    R090641_chip1.jpg

  • right side of chip 1 from lbcr.20100605.083801, I-Bessel:
    R083801_chip1.jpg
I Attachment Action Size Date Who Comment
R083801_chip1.jpgjpg R083801_chip1.jpg manage 77 K 21 Jun 2010 - 16:20 OlgaKuhn right side of chip 1 from lbcr.20100605.083801, I-Bessel
R090641_chip1.jpgjpg R090641_chip1.jpg manage 121 K 21 Jun 2010 - 16:13 OlgaKuhn right side of chip 1 from lbcr.20100618.090641, I-Bessel
compare_pupils.jpgjpg compare_pupils.jpg manage 46 K 18 Jun 2010 - 10:30 OlgaKuhn Comparison of pupils on 5 June and 18 June 2010.
Topic revision: r7 - 21 Jun 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback